OpenStudy (aye_my_peeps):

what is 4 2/7 x 2 1/5?

OpenStudy (aye_my_peeps):

what is \[4\frac{ 2 }{ 7 }\] times \[2\frac{ 1 }{ 5 }\]

OpenStudy (aye_my_peeps):

@Awolflover1 can u help me if your not doin anything

OpenStudy (aye_my_peeps):

or anyone

OpenStudy (stuck-help):

do you have answer choises

OpenStudy (aye_my_peeps):

no that is wat makes it harder

OpenStudy (anonymous):

9.42857142857

OpenStudy (stuck-help):

ok change these in to decimal form 1 divided by 5 and 2 divided by 7 @brejaye please no direct answers that breaks the rules of open study

OpenStudy (aye_my_peeps):

im sorry but im still confused

OpenStudy (anonymous):

change the 2/7 and the 1/5 to decimals so that you can multiply them

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (stuck-help):

i wasnt finished then add you number to the decimal of you fraction and then multiply them

OpenStudy (anonymous):

\[a+\frac{b}{c}=\frac{a*c+b}{c} \]

OpenStudy (aye_my_peeps):

ummm....

OpenStudy (anonymous):

type \[2\div7\] into your calculator

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

16/35 or in other words 0.4571428571

OpenStudy (anonymous):

I did it on a scientific calculator

OpenStudy (anonymous):

@Princeyboy4 let the person who asked the question learn how to solve it. So they know how to in the future

OpenStudy (sleepyjess):

I'm not sure if you got the answer or not, but first we need to convert these into an improper fraction. To do that, do what @robtobey said, \(\dfrac{a*c+b}c\). In the first fraction, we have 4 \(\dfrac 27\). a = 4 b = 2 c = 7 So now, we do \(\dfrac{4*7 + 2}7\). That gives \(\dfrac{30}7\). Now, we have 2 \(\dfrac 15\) for the second fraction. Do the same thing. a = 2 b = 1 c = 5 \(\dfrac{2*5 + 1}5\), that gives \(\dfrac{11}5\). Now, take those fractions and multiply across the top and bottom. \(\dfrac{30}7 * \dfrac{11}5\) 30*11 is the numerator of the answer 7*5 is the denominator of the answer. Hopefully that all makes sense :)

OpenStudy (anonymous):

@brejaye I'm just trying to help

OpenStudy (aye_my_peeps):

OMG THAT LOOKS CONFUSIN

OpenStudy (sleepyjess):

Just read through it and ask questions after if you still have any

OpenStudy (aye_my_peeps):

OK..

OpenStudy (anonymous):

@AYE_MY_PEEPS did you get it right

OpenStudy (aye_my_peeps):

no.. :(

OpenStudy (aye_my_peeps):

@Princeyboy4

OpenStudy (aye_my_peeps):

ok bye nvm thnks 4 tryin to help...

OpenStudy (stuck-help):